# Relevant or Random: Can LLMs Truly Perform Analogical Reasoning?

### v1

@@ Abstract
Analogical reasoning is a unique ability of humans to address unfamiliar challenges by transferring strategies from relevant past experiences. One key finding in psychology is that compared with irrelevant past experiences, recalling relevant ones can help humans better handle new tasks Ramachandran (2012). Coincidentally, the NLP community has also recently found that self-generating relevant examples in the context can help large language models (LLMs) better solve a given problem than hand-crafted prompts Yasunaga et al. (2024). However, it is yet not clear whether relevance is the key factor eliciting such capability, i.e., can LLMs benefit more from self-generated relevant examples than irrelevant ones? In this work, we systematically explore whether LLMs can truly perform analogical reasoning on a diverse set of reasoning tasks. With extensive experiments and analysis, we show that self-generated random examples can surprisingly achieve comparable or even better performance, e.g., 4 performance boost on GSM8K with random biological examples. We find that the accuracy of self-generated examples is the key factor and subsequently design two improved methods with significantly reduced inference costs. Overall, we aim to advance a deeper understanding of LLM analogical reasoning and hope this work stimulates further research in the design of self-generated contexts.
@@ 1 Introduction
A hallmark of human intelligence is that they can solve novel problems by drawing analogy from relevant past experiences, a concept known as analogical reasoning in cognitive science (Vosniadou and Ortony, 1989). As indicated by the name, recalling previously acquired relevant experiences can facilitate humans to better tackle new tasks, whereas irrelevant ones are rarely beneficial and can even be distracting Ramachandran (2012). For instance, when faced with a novel math problem about determinants (e.g., calculating the value of a given fourth-order determinant), humans can resolve this by reflecting upon the methodology employed to ascertain the value of a third-order determinant, whereas biological knowledge (e.g., how the human body regulates its temperature) can generally be considered irrelevant.
With the recent advancements in scaling up model size and data, LLMs have demonstrated impressive zero-shot and few-shot performance across various reasoning tasks, especially, through advanced prompting methods like chain-of-thought or CoT (Wei et al., 2022). Compared to common approaches such as zero or few-shot CoT (Zhou et al., 2022; Kojima et al., 2022; Wang et al., 2022b), recent work claims that LLMs are able to self-generate examples relevant to the query as context to better solve new problems (Webb et al., 2023; Yasunaga et al., 2024); see Fig. 1 for an example. However, it remains unclear whether relevance is the key to eliciting such capability in LLMs.
@@
To answer these questions, we empirically analyze the analogical reasoning abilities of GPT-3.5 (turbo) and Llama 2-Chat (Touvron et al., 2023c). Surprisingly, experimental results show that prompting LLMs to self-generate random examples can achieve comparable or even better performance on certain tasks, indicating that LLMs cannot always perform analogical reasoning. As for Q2, we point out through controlled experiments that the key factor is the accuracy of self-generated examples. Informed by these findings, we design two approaches that can outperform existing methods with significantly reduced inference costs. Specifically, we ask LLMs to randomly generate a few problems and manually verify their correctness, then use this fixed set of problems as in-context learning demonstrations for all test samples. Consistent observations across different model types consolidate the conclusions. We summarize the major contributions of our work below:
@@
With extensive experiments and analysis, we demonstrate the effectiveness and limitations of different types of self-generated contexts. Our code base is available at <redacted>.
@@ 2 Related Work
This work mainly explores whether LLMs can truly perform analogical reasoning. In light of this, we review three lines of research that form the basis of this work: large language models, chain-of-thought prompting, and analogical prompting.
2.1 Large Language Models
Ever since Brown et al. (2020) demonstrated that LLMs can achieve impressive few-shot and zero-shot performance on a variety of tasks, numerous efforts have been made to advance the development of LLMs. One line of research explored the benefits of scaling up LLMs, including Gopher (Rae et al., 2021), LaMDA (Thoppilan et al., 2022), MT-NLG (Smith et al., 2022) and PaLM (Chowdhery et al., 2022). Rather than scaling up model size alone, another line of research focused on training smaller models with better strategies or objectives, such as FLAN (Wei et al., 2021), T0 (Sanh et al., 2021), Chinchilla (Hoffmann et al., 2022), UL2 (Tay et al., 2022) and Flan-T5 (Chung et al., 2022).
The emergence of OpenAI’s ChatGPT/GPT-4 has greatly changed the research focus of the NLP community (Qin et al., 2023b; OpenAI, 2023). As reinforcement learning from human feedback (RLHF) contributed significantly to the success of these models (Schulman et al., 2017; Ouyang et al., 2022), researchers have explored various methods to improve RLHF (Yuan et al., 2023; Rafailov et al., 2023; Lee et al., 2023; Azar et al., 2023; Liu et al., 2024). To promote research of open-source LLMs, Meta released Llama series models (Touvron et al., 2023a, b), which serve as the solid foundation for a large number of open-source models (Taori et al., 2023; Chiang et al., 2023; Xu et al., 2023; Zhou et al., 2023). Furthermore, numerous endeavors have been dedicated to training powerful LLMs from scratch, instead of developing LLMs based on Llama (Li et al., 2023; Jiang et al., 2023; Bi et al., 2024; Mesnard et al., 2024).
@@ 2.2 Chain-of-Thought Prompting
Chain-of-thought (CoT) prompting induces LLMs to generate intermediate reasoning steps before generating the final answer (Wei et al., 2022), greatly improving the reasoning capabilities of LLMs. Typical CoT prompting approaches include few-shot CoT (Wei et al., 2022; Zhou et al., 2022; Wang et al., 2022b; Li et al., 2022; Wang et al., 2022a), taking several labeled demonstrations of the reasoning process, and zero-shot CoT, comprising only instructions like “Let’s think step by step” (Kojima et al., 2022; Zelikman et al., 2022; Zhang et al., 2023). Other ongoing research on CoT has also explored (i) optimizing the demonstration selection (Fu et al., 2022; Lu et al., 2022; Qin et al., 2023a), (ii) optimizing the quality of reasoning chains (Khot et al., 2022; Chen et al., 2022; Zhao et al., 2023; Shinn et al., 2023), and (iii) CoT in smaller language models (Magister et al., 2022; Ho et al., 2022; Fu et al., 2023).
2.3 Analogical Prompting
While few-shot CoT can provide more detailed reasoning guidance, it requires labeled examples which can be unavailable for a new task. To tackle this problem, Yasunaga et al. (2024) propose analogical prompting to guide LLMs to self-generate relevant exemplars as few-shot demonstrations, which is similar to analogical reasoning, i.e., humans can address new problems by drawing analogy from relevant past experience (Vosniadou and Ortony, 1989; Holyoak, 2012). In this work, we step forward to explore the intrinsic principle of analogical reasoning in LLMs. Specifically, we aim to investigate whether LLMs can authentically exhibit such reasoning capabilities and determine the extent to which the relevance of self-generated examples contributes to enhancing this process.
@@ 3 Methodology
In this work, we make minimal changes to the original prompt to guide LLMs to generate different types of irrelevant examples as context; see Fig. 2 and Fig. 3 for example prompts:
@@
Same as the original one, all newly designed prompts allow LLMs to generate irrelevant problems and solve the initial target problem in one pass. In addition, we show prompts for different methods on all datasets in Section A.1.
@@ 4.2 Main Results
As observed from Table 4, relevant examples are much more semantically similar to the query than irrelevant ones and the relevance score of ‘relevant’ is more biased towards ‘oracle’ rather than ‘random’ or ‘N/A’, demonstrating that LLMs indeed follow instructions to generate specific types of demonstrations.
@@
We refer to these two methods as and , and conduct experiments with them on GSM8K and MATH (see Fig. 5 for example prompts and outputs for generating math problems). Detailed prompts and outputs for different methods are provided in Section A.3. Following the original setting, we ask the LLM to randomly generate 5 examples for GSM8K and 3 examples for MATH. As observed from Table 8, ensuring the accuracy of self-generated examples does lead to better performance regardless of the problem type. Moreover, both ICL variants only need to generate examples once, which significantly reduces the inference cost and further demonstrates their superiority.
@@ 4.3 Further Analysis
Our experiments and analysis so far used GPT-3.5 as the LLM, which is closed-source and gets updated over time. To verify whether the observations and conclusions are consistent across different models and additionally for reproducibility, we extend the experiments to Llama-2-Chat (Touvron et al., 2023c). Specifically, we use vLLM (Kwon et al., 2023) to serve a Llama-2-70b-Chat model for experiments and report the results of different methods/variants on GSM8K in Table 9 and Table 10. We can draw similar observations: (i) self-generated relevant examples underperform all types of irrelevant ones, (ii) ‘GPT4-Calibration’ consistently outperforms the other two variants, and (iii) and perform better than other approaches, demonstrating that the conclusions can be generalized to different models.
@@ 5 Conclusion
This work has several limitations. First, due to the inference cost of ChatGPT111experiments done between 12/2023 and 04/2024, we conduct experiments on subsets of the test data for mathematical reasoning tasks. Besides, we include 3 datasets requiring different reasoning capabilities in this work. A further improvement could be to explore more diverse types of tasks.

### v2

@@ Abstract
Analogical reasoning is a unique ability of humans to address unfamiliar challenges by transferring strategies from relevant past experiences. One key finding in psychology is that compared with irrelevant past experiences, recalling relevant ones can help humans better handle new tasks Gentner and Smith (2012). Coincidentally, the NLP community has also recently found that self-generating relevant examples in the context can help large language models (LLMs) better solve a given problem than hand-crafted prompts Yasunaga et al. (2024). However, it is yet not clear whether relevance is the key factor eliciting such capability, i.e., can LLMs benefit more from self-generated relevant examples than irrelevant ones? In this work, we systematically explore whether LLMs can truly perform analogical reasoning on a diverse set of reasoning tasks. With extensive experiments and analysis, we show that self-generated random examples can surprisingly achieve comparable or even better performance, e.g., 4 performance boost on GSM8K with random biological examples. We find that the accuracy of self-generated examples is the key factor and subsequently design two improved methods with significantly reduced inference costs. Overall, we aim to advance a deeper understanding of LLM analogical reasoning and hope this work stimulates further research in the design of self-generated contexts.
@@ 1 Introduction
A hallmark of human intelligence is that they can solve novel problems by drawing analogy from relevant past experiences, a concept known as analogical reasoning in cognitive science (Vosniadou and Ortony, 1989). As indicated by the name, recalling previously acquired relevant experiences can facilitate humans to better tackle new tasks, whereas irrelevant ones are rarely beneficial and can even be distracting Gentner and Smith (2012). For instance, when faced with a novel math problem about determinants (e.g., calculating the value of a given fourth-order determinant), humans can resolve this by reflecting upon the methodology employed to ascertain the value of a third-order determinant, whereas biological knowledge (e.g., how the human body regulates its temperature) can generally be considered irrelevant.
With the recent advancements in scaling up model size and data, LLMs have demonstrated impressive zero-shot and few-shot performance across various reasoning tasks, especially, through advanced prompting methods like chain-of-thought (CoT) (Wei et al., 2022). Compared to common approaches such as zero or few-shot CoT (Zhou et al., 2022; Kojima et al., 2022; Wang et al., 2022b), Yasunaga et al. (2024) introduce LLM analogical reasoning, i.e., LLMs self-generate examples relevant to the query as context to better solve new problems; see Fig. 1 for an example. However, it remains unclear whether relevance is the key to eliciting such capability in LLMs. While several studies explore the influence of the relevance of demonstrations in in-context learning and CoT (Liu et al., 2022; Kim et al., 2022; Lyu et al., 2023; Chen et al., 2023; Yang et al., 2023; Wang et al., 2023; Yasunaga et al., 2024), none of them investigate whether self-generated relevant examples consistently outperform irrelevant ones in LLM analogical reasoning.
@@
To answer these questions, we empirically analyze the analogical reasoning abilities of GPT-3.5 (turbo) and Llama series models (Touvron et al., 2023). Surprisingly, experimental results show that prompting LLMs to self-generate random examples can achieve comparable or even better performance on certain tasks which is not in line with the key claim of analogical reasoning in Gentner and Smith (2012), indicating that LLMs cannot always perform analogical reasoning. As for Q2, we point out through controlled experiments that the key factor is the accuracy of self-generated examples. Informed by these findings, we design two approaches that can outperform existing methods with significantly reduced inference costs. Specifically, we ask LLMs to randomly generate a few problems and manually verify their correctness, then use this fixed set of problems as in-context learning demonstrations for all test samples. Consistent observations across different model types consolidate the conclusions. We summarize the major contributions of our work below:
@@
With extensive experiments and analysis, we demonstrate the effectiveness and limitations of different types of self-generated contexts.
@@ 2 Related Work
This work mainly explores whether LLMs can truly perform analogical reasoning. In light of this, we review two lines of research that form the basis of this work: chain-of-thought prompting and LLM analogical reasoning.
@@ 2.1 Chain-of-Thought Prompting
Chain-of-thought (CoT) prompting induces LLMs to generate intermediate reasoning steps before generating the final answer (Wei et al., 2022), greatly improving the reasoning capabilities of LLMs. Typical CoT prompting approaches include few-shot CoT (Wei et al., 2022; Zhou et al., 2022; Wang et al., 2022b; Li et al., 2022; Wang et al., 2022a), taking several labeled demonstrations of the reasoning process, and zero-shot CoT, comprising only instructions like “Let’s think step by step” (Kojima et al., 2022; Zelikman et al., 2022; Zhang et al., 2023). Other ongoing research on CoT has also explored (i) optimizing the demonstration selection (Fu et al., 2022; Lu et al., 2022; Qin et al., 2023), (ii) optimizing the quality of reasoning chains (Khot et al., 2022; Chen et al., 2022; Zhao et al., 2023; Shinn et al., 2023), and (iii) CoT in smaller language models (Magister et al., 2022; Ho et al., 2022; Fu et al., 2023).
2.2 LLM Analogical Reasoning
While few-shot CoT can provide more detailed reasoning guidance, it requires labeled examples which can be unavailable for a new task. To tackle this problem, Yasunaga et al. (2024) propose analogical prompting to guide LLMs to self-generate relevant exemplars as few-shot demonstrations, which is similar to analogical reasoning, i.e., humans can address new problems by drawing analogy from relevant past experience (Vosniadou and Ortony, 1989; Holyoak, 2012). In this work, we step forward to explore the intrinsic principle of LLM analogical reasoning. Specifically, we aim to investigate whether LLMs can authentically exhibit such reasoning capabilities and determine the extent to which the relevance of self-generated examples contributes to enhancing this process.
@@ 3 Methodology
In this work, we make minimal changes to the original prompt to guide LLMs to generate different types of irrelevant examples as context; see Fig. 2 for example prompts:
@@
Yasunaga et al. (2024) demonstrate that self-generating relevant examples can consistently outperform zero-shot CoT and few-shot CoT (hand-crafted examples or retrieved top- most similar training samples) on different tasks. Therefore, we do not include these two methods in our work. Interested readers can refer to the corresponding results and analysis in Yasunaga et al. (2024). In addition, we show prompts for different methods on all datasets in Section A.1.
@@ 4.2 Main Results
As observed from Table 4, relevant examples are much more semantically similar to the query than irrelevant ones and the relevance score of ‘relevant’ is more biased towards ‘oracle’ rather than ‘random’ or ‘N/A’, demonstrating that LLMs indeed follow instructions to generate specific types of demonstrations. Furthermore, we calculate the average similarity score between self-generated relevant examples and queries for BBH (0.46), which is slightly lower than the score of mathematical reasoning tasks (0.48). This result demonstrates that the difference in analogical reasoning performance between BBH (Table 1) and mathematical reasoning (Table 2) is not because LLMs can generate more relevant examples for BBH.
@@
It is worthwhile to note that while several papers explore how the correctness of demonstration answers influences in-context learning (Min et al., 2022; Yoo et al., 2022; Wei et al., 2023; Pan et al., 2023; Kossen et al., 2024), our work differs from them in the following aspects: (i) The examples in our work are generated by LLMs rather than real data from NLP benchmarks, i.e., randomly sampled from the training set. In addition, there are rationales (CoT) in self-generated examples, which are different from the input-label format of in-context learning investigated in these papers; and (ii) These studies mainly evaluate in-context learning on different classification or multi-choice datasets, i.e., the output space is a finite set. In contrast, we are evaluating mathematical reasoning tasks, where the output space is infinite.
Given the above findings, a natural question is:
[…]
We refer to these two methods as and , and conduct experiments with them on GSM8K and MATH (see Fig. 4 for example prompts and outputs for generating math problems). Detailed prompts and outputs for different methods are provided in Section A.3. Following the original setting, we ask the LLM to randomly generate 5 examples for GSM8K and 3 examples for MATH. As observed from Table 8, ensuring the accuracy of self-generated examples does lead to better performance regardless of the problem type. and achieve similar average performance, once again demonstrating that relevance does not matter (see Section A.4 for more analysis on relevance). Moreover, both ICL variants only need to generate examples once, which significantly reduces the inference cost and further demonstrates their superiority.
| Method | Task | ||
| GSM8K | MATH | Average | |
| 75.7 | 36.8 | 56.3 | |
| 73.8 | 36.2 | 55.0 | |
| Relevant | N/A | |||
| 71.2 | 74.9 | 75.3 | 75.9 | 74.3 |
@@ 4.3 Further Analysis
Robustness to Prompt Formats
To verify the robustness of different methods to prompt formats, we experiment with a new prompt paraphrased from the original one by GPT-4 and present the results on GSM8K in Table 10. We also observe better performance with irrelevant examples than relevant ones, showing the robustness.
@@
Our experiments and analysis so far used GPT-3.5 as the LLM, which is closed-source and gets updated over time. To verify whether the observations and conclusions are consistent across different models and additionally for reproducibility, we extend the experiments to Llama-2-Chat (Touvron et al., 2023). Specifically, we use vLLM (Kwon et al., 2023) to serve a Llama-2-70b-Chat model for experiments and report the results of different methods/variants on GSM8K in Table 11 and Table 12. We can draw similar observations: (i) self-generated relevant examples underperform all types of irrelevant ones, (ii) ‘GPT4-Calibration’ consistently outperforms the other two variants, and (iii) and perform better than other approaches, demonstrating that the conclusions can be generalized to different models.
We further conduct experiments with the latest Llama-3-8B-Instruct model. The results reported in Table 11 demonstrate the generalizability of the conclusions across different model scales. In addition, since investigating analogical reasoning requires LLMs to self-generate different types of problems, we only experiment with instruction-tuned LLMs to ensure that they can follow the given instructions.
Different Numbers of Demonstrations
@@ 5 Conclusion
This work has several limitations. First, due to the inference cost of ChatGPT111experiments done between 12/2023 and 05/2024, we conduct experiments on subsets of the test data for mathematical reasoning tasks. Besides, we include 3 datasets requiring different reasoning capabilities in this work. A further improvement could be to explore more diverse types of tasks.
@@ A.2 Detailed Results for Different Random Seeds
A.3 Prompts and Outputs for Example Generation
A.4 Guided Problem Generation
In addition to random problem generation in Section 4.2-Q2-1, we further investigate guided problem generation. Specifically, we randomly select 5 training samples to guide LLMs to self-generate relevant math problems. We then manually verify their correctness and use this fixed set of problems as ICL demonstrations for experiments. The performance of this approach (56.1) is slightly lower than that of (56.3), verifying that relevance is not the key influencing factor.